# A Multimodal Benchmark for Video Fake News Detection

### v1

@@ B.1.1 Details of Videos Category
B.2 Task definition
In order to assess the MLLMs with more balanced test data, we construct a new test-split, we call it category-test-split, and the dataset shown in Figure 20. Apart from perception sub-tasks ,we also define a few understanding and reasoning sub-tasks, the definition as following:
B.2.1 Perception
1) Shooting angles counting, SAC. More angles can record news from multi aspects, especially in accident scene footage. MLLMs were required to recognize and output the number of shooting angles after browsing the news video.
  2) Human identity recognition, HIR. Given a video, MLLMs were required to retrieve all identities that have appeared in the video. The task is multi-choice question with limited options.  3) Key Elements Temporal Grounding, KEG The Key elements temporal grounding task requires MLLMs to answer the time range in which the target elements appear in the video. The description of target elements is provided in the input, and the task is annotated by the process in the Figure 4. we use the Intersection over Union (IoU) metric, calculated as the ratio of the intersection to the union of predicted and labeled time intervals.
B.2.2 Understanding
1) News video type understanding, NTU. Video’s category depends on the theme of video, the task is a single-choice question, MLLMs’ are instructed to evaluate the theme of the news and output the type. All options can checked in Figure 20
  2) News video Theme understanding, NEU. In news video, almost all information is added and managed to reflect theme of the news video. Understanding theme of the video guide the detection to pay attention to the elements relevant to the key contents. MLLMs’ were guided to generate semantic description of video theme.
B.3 Metrics
To evaluate the capabilities of MLLMs across the 10 tasks, we define multiple metrics tailored to the task type and output format.
  1) Avg EHR(element hit rate): The metric measures MLLMs’ element hit rate in KEP task. Prediction is $\hat{y}=\{\hat{e_{0}},\hat{e_{1}}...\hat{e_{n}}\}$, corresponding label is $y=\{e_{0},e_{1}...e_{m}\}$. n is the number of elements in prediction, m is the number of elements in label. To ensure the highest possible accuracy of the results, each label element $e_{i}$ is evaluated by the GPT-4. The result of matching $r_{i}=M(e_{i},\hat{y}),r_{i}\in\{0,1\}$, M is GPT-4. $EHR_{j}=\sum_{i=0}^{m}r_{i}/m,0\leq j<t$, t is number of video. $AvgEHR=\sum_{j=0}^{t}EHR_{j}/t$.
  2) Avg ENHR(entity hit rate): The metric measures MLLMs’ entity hit rate in FDR task. Reasoning output is $\hat{y}$, corresponding label is $y=\{e_{0},e_{1}...e_{m}\}$, m is the number of label entities. To ensure the highest possible accuracy of the results, each label entity $e_{i}$ is evaluated by the GPT-4. The result of matching $r_{i}=M(e_{i},\hat{y}),r_{i}\in\{0,1\}$, M is GPT-4. $ENHR_{j}=\sum_{i=0}^{m}r_{i}/m,0\leq j<t$, t is number of video. $AvgENHR=\sum_{j=0}^{t}ENHR_{j}/t$.
  3) Avg KHR(knowledge hit rate): The metric measures MLLMs’ Knowledge hit rate in FDR task. Reasoning output is $\hat{y}$, corresponding label is $y=\{e_{0},e_{1}...e_{m}\}$, m is the number of label knowledge. To ensure the highest possible accuracy of the results, each label knowledge $e_{i}$ is evaluated by the GPT-4. The result of matching $r_{i}=M(e_{i},\hat{y}),r_{i}\in\{0,1\}$, M is GPT-4. $KHR_{j}=\sum_{i=0}^{m}r_{i}/m,0\leq j<t$, t is number of video. $\text{Avg KHR}=\sum_{j=0}^{t}KHR_{j}/t$.
  4) Avg RHR(rationale hit rate): The metric measures MLLMs’ rationale hit rate in FDR task. Reasoning output is $\hat{y}$, corresponding label is $y=\{e_{0},e_{1}...e_{m}\}$, m is the number of label rationales. To ensure the highest possible accuracy of the results, each label entity $e_{i}$ is evaluated by the GPT-4. The result of matching $r_{i}=M(e_{i},\hat{y}),r_{i}\in\{0,1\}$, M is GPT-4. $RHR_{j}=\sum_{i=0}^{m}r_{i}/m,0\leq j<t$, t is number of video. $\text{Avg RHR}=\sum_{j=0}^{t}RHR_{j}/t$.
  5) Avg AD(Absolute Distance): The metric measures MLLMs’ performance in SAC task. $\hat{d}$ is the number of shooting angles predicted by MLLMs, $d$ is ground truth. Absolute distance of each video is $AD_{i}=|\hat{d}-d|,0\leq i<n,$n is the number of video. $AvgAD=\sum_{i=0}^{t}AD_{i}/n$   6) Avg IoU(Intersection over Union): The metric measures MLLMs’ temporal grounding capacity in KEG task. Prediction time range is $\hat{y}=\{\hat{t_{0}},\hat{t_{1}}...\hat{t_{n}}\}$, corresponding label time range is $y=\{t_{0},t_{1}...t_{m}\}$. n is the number of time range in prediction, m is the number of time range in label. $IoU_{i}(\hat{t_{0}},t_{0})=\frac{|\hat{t_{0}}\cap t_{0}|}{|\hat{t_{0}}\cup t_{0}|},0\leq i<m$, m is the number of video, $\text{Avg IoU}=\sum_{j=0}^{m}IoU_{j}/m$.
For tasks involving creator-added content recognition, position detection, and color perception, we measure performance using accuracy based on evaluation results generated by GPT-4o. For tasks such as counting shooting angles and reasoning on three metrics (importance, emotional level, and relevance), we calculate the average absolute distance between the candidate and correct answers. To assess theme understanding, author-added content, and video shooting reasoning, we employ three metrics: Factual Consistency (FC), Theme Relevance (TR), and Completeness (CO). These metrics capture the degree of hallucination, alignment with the theme, and comprehensiveness of the output, respectively. For tasks involving human identity recognition (multiple-choice) and video type classification (single-choice), we evaluate performance using selection accuracy. To assess the temporal grounding of key elements, we use the Intersection over Union (IoU) metric, calculated as the ratio of the intersection to the union of predicted and labeled time intervals. All scores generated by the language model are normalized to a scale of 0 to 5. For final detection, in addition to using accuracy to evaluate the binary output, six evaluation metrics are defined for the reasoning process: entity hit rate, rationale hit rate, knowledge hit rate, theme relevance, factual consistency, and ROUGE-L. Among these metrics, the first three fall into the category of exact match, while the latter three belong to semantic match. Specifically, entity hit rate and theme relevance reflect the degree of correlation between the reasoning process and the video content; rationale hit rate, knowledge hit rate, and factual consistency reflect the reasoning capabilities of MLLMs; and ROUGE-L measures the semantic quality of the reasoning output.
B.4 Main Results
Table 4: Evaluation Result of CoT on Fake News Detection
Model CAC-MVFND-CoT OSF-MVFND-CoT ALL-MVFND-CoT MVFND-CoT Instruct-Tuning
@@ B.4 Main Results
GPT-4o-mini 72.24 78.81 72.00 78.87 71.31 74.66 71.05 72.35 - -
Qwen2.5vl-72b 77.95 74.08 78.35 74.19 77.56 78.12 77.00 74.23 - -
Qwen2.5vl-32b 74.90 69.16 74.76 74.68 75.43 70.69 74.90 76.05 - -
@@
Qwen2.5-7b-Instruct 57.23 53.71 57.50 55.39 57.76 55.36 58.43 58.23 81.14 80.26

### v2

@@ A.1 Details of Videos Category
Figure 7: Details of Training Dataset.
Appendix B Task Details
As demonstrated in Figure 8, 9, 10. We define 10 tasks spanning perception, understanding, and reasoning, with diverse question formats including single selection, multiple selections, open-ended output, and structured output. In Key Elements Temporal Grounding (KEG), the model is instructed to identify the time range during which target elements appear, with each question containing at least one key element. Key Elements Perception (KEP) requires the model to retrieve key elements from video news and generate their semantic descriptions in a structured format. In Distinguish Creator-added Content and Original Shooting Footage (DCS), the model must recognize author-added content, such as on-screen text, and distinguish it from original shooting footage. Creator-added Text Color Perception (CCP) evaluates the model’s ability to perceive the font color of target text. For Creator-added Text 2D-position Perception (CPP), the model identifies the screen position of target text, with semantic descriptions of all potential positions provided in the prompt to ensure accurate assessment. Shooting Angles Counting (SAC) requires the model to detect transitions between shooting angles and count their occurrences. In Human Identity Recognition (HIR), the model identifies all persons appearing in the video news through a multiple selection format. News Video Type Understanding (NTU) tasks the model with comprehending overall content to determine the news type from provided options. News Video Theme Understanding (NEU) requires the model to describe the global topic of the news video, including events, locations, and temporal information. Finally, Final Detection Reasoning (FDR) instructs the model to synthesize perceived elements, understood content, and reasoning rationales to assess the veracity of the news video.
Figure 8: Definition of Evaluation Tasks
Figure 9: Definition of Evaluation Tasks
Figure 10: Definition of Evaluation Tasks. In the model output of the FDR process, the blue text represents entities identified in the video, including author-added text and original footage. The orange text denotes the rationales used during reasoning, comprising logical and empirical bases for determining authenticity. The green text indicates external knowledge related to the video employed during the reasoning process. These three components are used to evaluate the ENHR, KHR, and RHR of the reasoning process. The "Answer" is used to calculate the accuracy rate of the reasoning results.
Appendix C Data Statistics
We annotate 2,871 training videos, each accompanied by visual information, semantic descriptions, and reasoning evidence. As shown in Figure 7, the training split contains an average of 4.58 key elements per video, totaling 8,233 elements. Key elements have an average duration of 13.59 seconds, and the number of key shots exhibits diverse distributions. The dataset maintains balanced representation across 6 news categories.
Appendix D Rationale Validation
During reasoning, MLLMs leverage their background knowledge and visual information from videos to generate interpretable, evidence-based outputs. These outputs consist of multiple rationales that collectively support the final detection judgment (fake or real). To validate the factual correctness of these rationales, we conduct a verification study on the original FakeSV dataset, examining whether the visual discrepancies described in the rationales actually exist in the source videos. Beyond validation, we also provide practical explanations for identified discrepancies. We categorize rationales into two types: those relevant to creator-added content (CAC) and those relevant to original shooting footage (OSF). For CAC rationales validation, we employ the automated pipeline illustrated in Figure 11 to extract visual features. For OSF rationales validation, we leverage the expert annotations in POVFNDB for validation. Our empirical analysis reveals that the distinguishing characteristics between fake and real news described in MLLM-generated rationales are corroborated by statistical patterns observed in the original FakeSV dataset. This empirical validation enhances the reliability of the MLLM reasoning process.
D.1 CAC Features Extraction
Creator-added content (CAC) in news videos predominantly consists of textual overlays inserted using specialized editing software. These overlays serve various functions, including providing event context through captions, introducing key figures, and displaying distinctive graphical identifiers from authoritative media outlets. To enable quantitative analysis of CAC rationales, we develop an automated extraction pipeline as shown in Figure 11. First, we uniformly sample 16 frames per video using FFmpeg with temporally equidistant intervals. Second, we apply PaddleOCR to detect text regions in each frame, obtaining bounding boxes and region screenshots. Since PaddleOCR occasionally detects text from the original shooting footage (e.g., signs shown in Figure 14), we manually filter such instances to retain only creator-added text regions. Third, we employ Qwen2.5-VL-72B with designed prompt templates to extract semantic attributes of each text region, including content, color, 2D position, and aspect ratio. Finally, domain experts refine the MLLM-generated descriptions to ensure accuracy.
Figure 11: Creator-added Content Extraction Process
D.2 CAC Rationales Validation
Creator-added text is typically used in video news to highlight key information, such as sites, date. We select two rationales about text font color, screen position, emotion level are frequently cited by MLLMs. We utilize visual information generated by the pipeline in the Figure 11. We employee the PaddleOCR tools to extract text regions from video and process them with the pipeline in Figure 11.
Figure 12: Hue distribution in creator-added content
of fake/real news video
D.2.1 Color Distribution
Rationale-1. Yellow is a professional typographic color choice for video news content, offering high visibility and prominence while maintaining a neutral-to-bright aesthetic. This color selection minimizes emotional bias in viewers’ information processing and enhances the perceived authenticity of news content.
Rationale-2. Although red typography achieves high visual salience, it induces emotional arousal in audiences, which contradicts the factual nature of news dissemination and consequently diminishes perceived authenticity.
Analysis. Prior work has shown that font color affects news comprehension, with hue serving as the primary color identifier Zhou et al. (2022). To investigate whether color usage differs between real and fake news, we conduct a statistical analysis comparing pixel-level hue distributions in creator-added text across both categories. As shown in Figure 12, we observe significant divergence in hue distribution between real and fake news videos, particularly in red and yellow ranges. This finding aligns with MLLM-generated rationales that frequently cite font color as a veracity indicator.
We attribute this phenomenon to differences in production practices and intent. Fake news producers tend to employ emotionally salient colors (particularly red) to manipulate audience perception and enhance credibility of false claims. In contrast, creators of authentic news use formal, neutral color schemes to guide viewer attention to substantive information. Furthermore, the prevalence of high-saturation red in fake news may reflect rushed production workflows, where creators prioritize immediate visual impact over context-appropriate color selection. This production haste results in overreliance on attention-grabbing colors without consideration for content alignment.
(a) Distribution of text region in fake news video
(b) Distribution of text region in real news video
Figure 13: Comparison of heatmaps
D.2.2 Spatial Distribution
Rationale-1. The text describing the event location and time is positioned at the top-center of the screen, consistent with conventional placement patterns in real news videos.
Rationale-2. Text positioned at the screen center occludes the original footage, severely interfering with viewers’ access to authentic visual content. This layout pattern is commonly employed in fake news to obscure factual evidence, thereby reducing perceived authenticity.
Analysis. As illustrated in Figure 13, the spatial distribution of text regions exhibits notable differences between real and fake news videos, consistent with patterns described in MLLM-generated rationales. We attribute the observed distributional discrepancy to two factors.1) Fake news producers strategically overlay text on video footage to obscure factual details in the original visuals and redirect viewer attention toward fabricated textual narratives. 2) Fake news creators lack professional training in screen-text composition. Unlike authentic news production, where text placement follows established design principles that balance information salience with minimal occlusion of original content, fake news exhibits arbitrary positioning that disregards these compositional best practices.
Figure 14: CAC in green box,OSF in red box.
Figure 15: Individual with identity
D.3 Original Shooting Footage
Original shooting footage (OSF) typically provides first-hand documentation of events, particularly for breaking incidents. Unlike CAC, OSF is less susceptible to manipulation, thereby offering higher credibility. To investigate whether OSF-related features correlate with news veracity, we select rationales from MLLM reasoning that reference OSF characteristics and analyze their distribution across real and fake news videos using annotations from POVFNDB.
D.3.1 Key footage Distribution
Figure 16: The proportion of key footage appearing in different time segments of fake and real news videos. Here, "on_site" refers to footage from the event scene, "close_up" indicates close-up shots of individuals related to the event, and "declaration" represents statements from authoritative institutions regarding the event. "ts0-ts7" denotes segments of the video divided uniformly along the timeline, where "ts0" refers to the first segment of the video, "ts7" refers to the final segment, and so forth.
Rationale-1 The video concludes with screenshots of official police documents detailing the incident and involved parties, presenting authoritative information to viewers. This placement pattern—authoritative evidence at video end—is characteristic of authentic news reports and enhances perceived credibility.
Rationale-2 The video reports a breaking news event but lacks on-site footage, instead presenting multiple unrelated scenes. This absence of relevant visual evidence reduces perceived authenticity.
Analysis Key frames are closely tied to news themes and hold significant evidentiary value for veracity detection. In professional news production, critical information is typically positioned in opening segments to engage viewers, while supplementary content such as disclaimers and announcements appears in closing segments. To examine temporal distribution patterns, we analyze three types of salient visual content: event scenes, human faces, and textual statements. We uniformly partition each video into eight temporal segments and employ Qwen2.5-VL-72B to detect the presence of these three frame types in each segment. Figure 16 illustrates the resulting distributions, which align with patterns described in MLLM-generated rationales. We attribute the observed differences between real and fake news to two factors: fake news producers’ deliberate attempts to obscure factual evidence and their lack of professional training in news composition conventions.
D.3.2 Subject Identity Distribution
Figure 17: Distribution of subject identity in
fake(external)/real(internal) news video.
Rationale-1 The video features interviews with law enforcement personnel, with names displayed via on-screen captions. This practice is commonly employed in incident reporting to enhance news credibility.
Rationale-2 Despite being framed as breaking incident news, the video contains no appearances by event-related individuals, substantially reducing credibility.
Rationale-3 The video documents an incident and includes footage of the victim receiving medical treatment, with on-screen name identification. This verifiable content substantially enhances perceived news authenticity.
Analysis News videos frequently feature human subjects to convey key event information through specific visual formats, such as face-to-face interviews and press conference footage. These individuals include involved parties, victims, law enforcement personnel, and self-media creators, among others. However, the reliability of information provided by different identity types varies significantly Pereira et al. (2023). To investigate this phenomenon, we analyze the proportional distribution of identity types across entire video durations. Figure 17 reveals distinct distributional patterns between real and fake news. Perpetrators, victims, and official personnel appear more frequently in real news videos, with law enforcement officers appearing more than twice as often in real news compared to fake news. Notably, 22.6We attribute these disparities to two primary factors. First, real news producers possess institutional access to law enforcement personnel involved in reported events, whereas fake news creators lack such credentials. Second, the inclusion of authoritative figures serves as an essential verification mechanism to mitigate audience skepticism in authentic news production—an element that fake news producers deliberately avoid. Conversely, fake news videos exhibit higher proportions of both absent human subjects and self-media creators, reflecting their reliance on unofficial sources and reduced evidentiary standards.
Figure 18: Number of shooting angles in the fake and real video news
D.3.3 Relevant shooting angles
Rationale-1 The video captures the event from multiple perspectives, enabling viewers to comprehend the complete context and thereby enhancing perceived veracity.
Rationale-2 The video contains only a single static image of the event scene, preventing cross-verification of news content and reducing perceived authenticity.
Analysis We leverage annotated data from the benchmark to analyze the distribution of camera shots between fake and real news videos. As shown in Figure 18, fake news exhibits a higher proportion of videos with 0-1 shots, whereas real news dominates in videos with more than one shot. These patterns align with MLLM-generated rationales. To provide comprehensive and objective coverage of news events, professional creators employ multiple shooting angles within their videos. This approach enables audiences to gain a thorough understanding of event progression while enhancing the credibility and persuasiveness of news reports Bock et al. (2023). We attribute the observed disparities to three primary factors. First, fake news producers typically lack sufficient shooting materials about the reported events, limiting their ability to present multiple perspectives. Second, reduced shot diversity prevents audiences from fully comprehending the facts, which paradoxically serves the obfuscation goals of disinformation creators. Third, fake news producers often lack the professional videography skills required for seamless camera shot transitions in news production.
Appendix E Additional Experiments
We evaluated the performance of MLLMs across various capabilities on the POVFND task, utilizing both process-based and result-based tasks within the benchmark.We further verify the effectiveness of different visual prompts by comparing three variants: CAC-POVFND-CoT, OSF-POVFND-CoT, and ALL-POVFND-CoT, finding that OSF yields the most substantial performance gains. Through further analysis of the evaluation process, we identified several insights and discussed their underlying intrinsic causes.
Model CAC-POVFND-CoT OSF-POVFND-CoT ALL-POVFND-CoT POVFND-CoT Instruct-Tuning
@@ Appendix E Additional Experiments
GPT-4o-mini 72.24 73.56 72.00 73.45 71.31 74.66 71.05 72.35 - -
Qwen2.5VL-72b 77.95 74.08 78.35 74.19 77.56 78.12 77.00 76.05 - -
Qwen2.5VL-32b 74.90 69.16 74.76 74.68 75.43 70.69 74.90 74.23 - -
@@
Qwen2.5VL-7b-Instruct 57.23 53.71 57.50 55.39 57.76 55.36 58.43 58.23 81.14 80.26
Table 7: Evaluation results of different CoT frameworks on fake news detection. CAC-POVFND-CoT integrates creator-added content analysis (color, position, text), OSF-POVFND-CoT incorporates original shooting footage analysis (key elements, shooting angles, human identities), and ALL-POVFND-CoT combines both CAC and OSF components with CoT reasoning.